# Quantum Rationale-Aware Graph Contrastive Learning for Jet Discrimination

### v6

@@ 3.1.6 Enforcing Infrared and Collinear Safety
To ensure compliance with infrared and collinear (IRC) safety, we adopt perturbation and regularization techniques inspired by the principles of QCD, as outlined by Dillon et al. [26]. Further theoretical and implementation details are provided in Appendix C.2.1.
@@ 3.1.7 Data Splitting
To ensure manageable computational complexity and adapt the model for quantum processing, we focused on jets containing at least ten particles, resulting in a dataset of $N=1,997,445$ jets, of which $997,805$ were classified as quark jets. Unlike classical graph neural networks (GNNs), which offer flexibility in adjusting the number of hidden features, quantum networks are constrained by the scaling of quantum states and Hamiltonians. Specifically, the quantum computational cost scales as $2^{n}$, where $n$ represents the number of qubits, corresponding to the number of nodes ($n_{\alpha}$) in the graph. Each node represents a particle in the jet, making jets with many particles challenging to handle due to the exponential growth in quantum computational requirements.
To address the challenge of varying particle numbers in jets, we simplified the problem by limiting the number of active nodes (particles) per jet to $n_{\alpha}=7$. This was done by selecting the seven particles with the highest transverse momentum ($p_{T}$) from each jet. Consequently, each jet graph is represented by a feature set $h_{\alpha}=(h_{\alpha}^{(1)},h_{\alpha}^{(2)},\ldots,h_{\alpha}^{(7)})$, where each $h_{\alpha}^{(i)}\in\mathbb{R}^{8}$ corresponds to the enriched feature vector of a particle. The complete representation of each jet is thus given by $h_{\alpha}\in\mathbb{R}^{7\times 8}$, capturing key physical attributes of the selected particles. A subset of $N=12,500$ jets was randomly selected for model training, with 10,000 jets used for training, 1,250 for validation, and 1,250 for testing. These subsets maintained the original class distribution, resulting in 4,982 quark jets in the training set, 658 in the validation, and 583 in the testing set.
@@ 3.2.1 Quantum Rationale Generator (QRG)
The QRGCL model substitutes its classical RG (CRG) [7] with a quantum RG (QRG). This component is crucial in generating augmented graph representations by assigning significance scores to each node. The QRG is built using a 7-qubit parameterized quantum circuit (PQC), where each qubit represents a node in the graph.
@@
$H|0\rangle=\frac{1}{\sqrt{2}}(|0\rangle+|1\rangle).$ (2)
Figure 4: QRG circuit of the proposed QRGCL. The circuit operates on seven qubits, with each qubit corresponding to a node in the graph. The top portion shows the data encoding stage, where each qubit is initialized using an $H$ gate followed by RXRX-based angle encoding node features. CRZCRZ gates encode edge relationships between qubits. The bottom portion includes parameterized rotations (RXRX, RYRY, RZRZ) for adaptable representations and entanglement layers using SWAPSWAP gates. Measurement results are obtained on a computational basis, with classical registers collecting the output.
Next, node feature vectors are encoded using rotation gates, with various encoding options such as $RX$, RYRY, RZRZ, or $H$ gates. The specific choice of encoding can be customized, with $RX$ encoding being implemented for proposed angle-based representations: Node feature vectors $\mathbf{x}_{i}$ are encoded as rotation angles using $RX$, RYRY, or $RZ$ gates. For example, the $RX$ gate is defined as:
@@
where $\theta$ is determined by the value of $\mathbf{x}_{i}$. Edge relationships are encoded using controlled-phase ($CRZ$) gates between pairs of qubits. This is a diagonal, asymmetric gate that applies a phase to the target qubit depending on the control qubit’s state:
@@
$U3(\theta,\phi,\lambda)=\begin{pmatrix}\cos(\theta/2)&-e^{i\lambda}\sin(\theta/2)\\ e^{i\phi}\sin(\theta/2)&e^{i(\phi+\lambda)}\cos(\theta/2)\end{pmatrix}$ (5)
which allows the QRG to learn adaptable representations through trainable parameters $\theta$, $\phi$, and $\lambda$. These parameterized gates form the trainable part of the circuit, allowing the QRG to adaptively learn the significance scores based on the input data during training. The parameters of these gates are optimized through gradient-based approaches. Entanglement is introduced using a flexible choice of entanglement layers, such as $CNOT$, CZCZ, SWAPSWAP, or their butterfly-pattern variants. These entanglement patterns are designed based on the graph’s structure and ensure that important correlations between nodes are captured. For example, $SWAP$ gates can exchange quantum states between qubits, preserving relationships between specific nodes:
$SWAP=\begin{pmatrix}1&0&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&1\end{pmatrix}$. (6)
The output of the QRG is obtained by measuring the qubits on a computational basis. The squared amplitudes of states with a Hamming weight of 1 (e.g., $|00001\rangle,|00010\rangle,\ldots$) provide node significance scores normalized into a discrete probability distribution. With the QRG generating significance scores, augmented views are created for the downstream CL process. For an optimistic view, nodes with the highest significance scores are retained, preserving the most relevant information for classification. Negative views, on the other hand, are constructed by using less significant nodes or by introducing random variations, helping the model learn to distinguish between genuinely informative structures and noise.
@@ 3.2.2 Encoder Network
We used the ParticleNet [2] model as our encoder to convert the augmented views of input particle features into low-dimensional embeddings. ParticleNet is a graph-based neural network optimized for jet tagging, leveraging dynamic graph convolutional neural networks (DGCNN) to process unordered sets of particles, treating jets as particle clouds. More details about ParticleNet are shown in Appendix D.1.
@@ 3.2.3 Projection Head
Our architecture incorporates a projection head that effectively maps high-dimensional embeddings from the encoder into a compact, CL-optimized latent space. This distinction allows the encoder to learn features suited for downstream tasks while the projection head focuses on maximizing contrastive effectiveness. This projection head employs an initial linear transformation to reduce the dimensionality of the input embedding (128-dimensional) to a latent space of the same dimensionality, followed by a ReLU activation and a final linear layer. This layered design allows the projection head to align representations tightly with the CL objective by enforcing mutual information between anchor and rationale representations in the latent space.
@@ 3.2.4 Quantum-Enhanced Contrastive Loss
QRGCL model uses a carefully designed loss function that integrates multiple elements: InfoNCE, alignment, uniformity, rationale-aware loss (RA loss), and contrastive pair loss (CP loss), to optimize the learning of quantum-enhanced embeddings. More details about these losses are shown in Appendix D.2.
@@ 3.2.5 Classifier Head
The classifier head follows our GCL framework and consists of a simple neural network with a 128-neuron single linear layer and is designed to perform binary jet discrimination, mapping features to a single output with a sigmoid activation for probabilistic classification.
@@ 3.3 Benchmark Models
We developed two classical models (GNN and EQGNN), three quantum models (QGNN, EQGNN, and QCL), five hybrid classical-quantum models (CQCL, 3 QCGCL variants, and QRGCL with $RX+H$ encoding) (see Table 3), and the classical counterpart of QRGCL, i.e., CRGCL (see Table 2). More details are shown in Appendix E.
Table 1: Hyperparameter optimization results of the proposed QRGCL. Bold indicates the best performance.
Parameter type Parameter Test Accuracy ($\uparrow$) AUC ($\uparrow$) F1 Score ($\uparrow$)
Encoder type Amplitude 69.30% 74.69% 69.30%
IQP 68.90% 76.10% 68.47%
Displacement-amplitude 66.17% 72.10% 65.97%
Displacement-phase 70.17% 76.22% 69.13%
RY 67.00% 73.93% 67.00%
RX 70.80% 76.30% 69.60%
RZ 62.90% 67.32% 62.80%
H 70.83% 76.71% 69.65%
Phase 68.50% 74.21% 68.30%
Learning rate 1E-03 70.30% 76.20% 70.29%
1E-04 68.60% 74.41% 68.39%
1E-02 69.50% 75.81% 69.47%
3E-03 63.90% 68.83% 63.70%
5E-03 66.00% 70.89% 65.99%
Entanglement type CNOT 69.80% 76.58% 69.79%
CZ 68.20% 73.45% 68.14%
SWAP 71.00% 76.20% 70.79%
CNOT Butterfly 68.00% 74.63% 67.98%
CZ Butterfly 67.70% 73.20% 67.69%
SWAP Butterfly 68.00% 73.21% 67.79%
[…]
Table 2: Comparison between classical and quantum RG of RGCL. Bold indicates the best performance.
@@
Table 3: Performance benchmarking and ablation test of the proposed QRGCL. Bold indicates the best performance.
Model Test accuracy ($\uparrow$) AUC ($\uparrow$) F1 score ($\uparrow$) # Parameters ($\downarrow$) $n_{\alpha}$ ($\downarrow$) $n_{layer}$ ($\downarrow$) Batch size Epoch Encoder

### v7

@@ 3.1.6 Enforcing Infrared and Collinear Safety
To ensure compliance with infrared and collinear (IRC) safety, we adopt perturbation and regularization techniques inspired by the principles of QCD, as outlined by Dillon et al. (dillon_symmetries_2022). Further theoretical and implementation details are provided in Appendix A.4.1.
@@ 3.1.7 Data Splitting
To ensure manageable computational complexity and adapt the model for quantum processing, we focused on jets containing at least ten particles, resulting in a dataset of $N=1,997,445$ jets, of which $997,805$ were classified as quark jets. Unlike classical GNNs, which offer flexibility in adjusting the number of hidden features, quantum networks are constrained by the scaling of quantum states and Hamiltonians. Specifically, the computational cost of simulating the quantum state vector on classical hardware scales as $2^{n}$, where $n$ represents the number of qubits, corresponding to the number of nodes ($n_{\alpha}$) in the graph. Each node represents a particle in the jet, making jets with many particles challenging to handle due to the exponential growth in quantum computational requirements.
To address the challenge of varying particle numbers in jets, we simplified the problem by limiting the number of active nodes (particles) per jet to $n_{\alpha}=7$. This truncation to $n_{\alpha}=7$ is a constraint imposed by the exponential cost of classically simulating quantum state vectors ($2^{N}$). While this selection acts as a groomer that may remove soft radiation characteristic of gluon jets, selecting the highest transverse momentum ($p_{T}$) constituents ensures adherence to IRC safety by retaining the dominant kinematic energy flow. This was done by selecting the 7 particles with the highest $p_{T}$ from each jet. Consequently, each jet graph is represented by a feature set $h_{\alpha}=(h_{\alpha}^{(1)},h_{\alpha}^{(2)},\ldots,h_{\alpha}^{(7)})$, where each $h_{\alpha}^{(i)}\in\mathbb{R}^{8}$ corresponds to the enriched feature vector of a particle. The complete representation of each jet is thus given by $h_{\alpha}\in\mathbb{R}^{7\times 8}$, capturing key physical attributes of the selected particles. A subset of $N=12,500$ jets was randomly selected for model training, with 10,000 jets used for training, 1,250 for validation, and 1,250 for testing. These subsets maintained the original class distribution, resulting in 4,982 quark jets in the training set, 658 in the validation, and 583 in the testing set.
Figure 3: Overview of the proposed QRGCL framework. Given an input jet represented as a graph (anchor graph), a QRG assigns importance scores to nodes (particles) using a parameterized quantum circuit. Based on these scores, a rationale subgraph and a corresponding complement subgraph are sampled to construct multiple rationale-aware augmented views. These views are processed by a shared ParticleNet GNN encoder, producing graph-level embeddings via global average pooling. The embeddings are passed through a projection head to obtain representations used for CL. The framework jointly optimizes the QRG, encoder, and projection head by minimizing a combined objective that includes rationale-aware, alignment, uniformity, contrastive-pair, and InfoNCE losses. During fine-tuning, the learned representations are fed into a lightweight classifier head for supervised discrimination of quark–gluon jets.
@@ 3.2.1 Quantum Rationale Generator (QRG)
The QRGCL model substitutes its classical RG (CRG) (li2022) with a quantum RG (QRG). This component is crucial in generating augmented graph representations by assigning significance scores to each node. The QRG is built using a 7-qubit parameterized quantum circuit (PQC), where each qubit represents a node in the graph.
Permutation symmetry.
Our QRG processes the $n_{\alpha}=7$ selected constituents in a fixed order (sorted by $p_{T}$), hence the rationale-scoring subroutine is not permutation-equivariant. However, the downstream ParticleNet encoder operates on unordered particle sets/graphs with symmetric neighborhood aggregation (EdgeConv) and pooling, preserving permutation invariance of the learned graph-level representation.
@@
$H|0\rangle=\frac{1}{\sqrt{2}}(|0\rangle+|1\rangle)$ (2)
Figure 4: QRG circuit of the proposed QRGCL. The circuit operates on seven qubits, with each qubit corresponding to a node in the graph. Symbols denote: $H$ (Hadamard gate) for superposition; $R_{X},R_{Y},R_{Z}$ (Rotation gates) for feature encoding; CRZCRZ (Controlled-Rotation $Z$) for entangling edge topology; and the meter symbols for measurement in the computational basis. The top portion shows the data encoding stage, where each qubit is initialized using an $H$ gate followed by RXRX-based angle encoding node features. CRZCRZ gates encode edge relationships between qubits. The bottom portion includes parameterized rotations (RXRX, RYRY, RZRZ) for adaptable representations and entanglement layers using SWAPSWAP gates. Measurement results are obtained on a computational basis, with classical registers collecting the output.
Next, the classical node feature vectors are embedded into the quantum state using parameterized rotation gates (e.g., $R_{X},R_{Y},R_{Z}$) or Hadamard-based encodings, mapping classical data to the Hilbert space of the quantum circuit. The specific choice of encoding can be customized, with $RX$ encoding being implemented for proposed angle-based representations. Node feature vectors $\mathbf{x}_{i}$ are encoded as rotation angles using $RX$, RYRY, or $RZ$ gates. For example, the $RX$ gate is defined as:
@@
[…]
Edge relationships are encoded using controlled-phase ($CRZ$) gates between pairs of qubits. The matrix representation is defined in the standard computational basis $\{|00\rangle,|01\rangle,|10\rangle,|11\rangle\}$, where the first qubit corresponds to the control state and the second to the target. This is a diagonal, asymmetric gate that applies a phase to the target qubit depending on the control qubit’s state.
@@
$U3(\alpha,\beta,\gamma)=\begin{pmatrix}\cos(\alpha/2)&-e^{i\gamma}\sin(\alpha/2)\\ e^{i\beta}\sin(\alpha/2)&e^{i(\beta+\gamma)}\cos(\alpha/2)\end{pmatrix}$ (5)
which allows the QRG to learn adaptable representations through trainable parameters $\alpha$, $\beta$, and $\gamma$. These parameterized gates form the trainable part of the circuit, allowing the QRG to adaptively learn the significance scores based on the input data during training. The parameters of these gates are optimized through the Adam optimizer.
The entanglement layers utilize a fixed circular topology, where the $i$-th qubit acts as the control for the $(i+1)$-th target qubit (with the last qubit controlling the first), ensuring complete pairwise correlation across the circuit layers. Entanglement is introduced using a fixed topology of two-qubit gates (specifically $SWAP$ gates in a circular pattern) to capture correlations between the sequence of input particles. These entanglement patterns are designed based on the graph’s structure, ensuring that important correlations between nodes are captured. For example, $SWAP$ gates can exchange quantum states between qubits, preserving relationships between specific nodes:
$SWAP=\begin{pmatrix}1&0&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&1\end{pmatrix}$ (6)
The output of the QRG is obtained by measuring the qubits on a computational basis. The squared amplitudes of states with a Hamming weight of 1 (e.g., $|0000001\rangle,|0000010\rangle,\ldots$) provide node significance scores normalized into a discrete probability distribution. To ensure the framework remains end-to-end differentiable, we utilize the Gumbel-Softmax reparameterization trick (jang2017categorical). This allows gradients to flow from the contrastive loss function back through the discrete sampling of rationale subgraphs, updating the parameters of the QRG and the projection layer. With the QRG generating significance scores, augmented views are created for the downstream CL process. For an optimistic view, nodes with the highest significance scores are retained, preserving the most relevant information for classification. Negative views, on the other hand, are constructed by using less significant nodes or by introducing random variations, helping the model learn to distinguish between genuinely informative structures and noise.
@@ 3.2.2 Encoder Network
We used the ParticleNet (qu_jet_2020) model as our encoder to convert the augmented views of input particle features into low-dimensional embeddings. ParticleNet is a graph-based neural network optimized for jet tagging, leveraging dynamic graph convolutional neural networks (DGCNN) to process unordered sets of particles, treating jets as particle clouds. The encoder is initialized with random weights and is fully trainable during the contrastive pre-training phase, allowing it to learn optimal representations alongside the rationale generator. It is only during the downstream supervised fine-tuning that the encoder weights are frozen. More details about ParticleNet are shown in Appendix B.1.
@@ 3.2.3 Projection Head
Our architecture includes a projection head that maps the 128-dimensional encoder output to a 128-dimensional latent space optimized for the contrastive loss function. Following the design principles of SimCLR (pmlr-v119-chen20j), we preserve the encoder output dimensionality (128) rather than compressing it directly, as a non-linear projection head has been shown to mitigate information loss prior to applying the contrastive objective. The projection head consists of a two-layer MLP with an intermediate ReLU activation, which transforms the encoder representations while maintaining the same latent dimensionality. CL is performed on these projected embeddings by maximizing the InfoNCE objective, which implicitly enforces mutual information between anchor and rationale representations. This separation enables the encoder to learn features transferable to downstream tasks, while the projection head specializes in aligning representations for effective contrastive optimization.
@@ 3.2.4 Quantum-Enhanced Contrastive Loss
QRGCL model uses a carefully designed loss function that integrates multiple elements: InfoNCE, alignment, uniformity, rationale-aware loss (RA loss), and contrastive pair loss (CP loss), to optimize the learning of quantum-enhanced embeddings. More details about these losses are shown in Appendix B.2.
@@ 3.2.5 Classifier Head
The classifier head, employed during the supervised fine-tuning phase, consists of a 128-neuron single linear layer ($d_{model}\to 1$) followed by a sigmoid activation. This simple architecture is designed to map the frozen representations to a probability score for binary jet discrimination, ensuring that performance gains are attributable to the quality of the learned embeddings rather than the complexity of the readout layer.
@@ 3.3 Benchmark Models
To validate the effectiveness of QRGCL, we benchmark against a diverse set of architectures categorized into Classical, Quantum, and Hybrid models (see Table 1 and Appendix C for detailed specifications):
Classical Baselines.
We employ a standard GNN (EdgeConv-based) and an Equivariant GNN (EGNN) to evaluate the impact of geometric symmetry preservation in the classical domain. We also compare against CRGCL (Table 2), the classical counterpart of our proposed method, to isolate the quantum advantage.
Quantum Baselines.
We utilize QGNN and EQGNN, which map graph structures directly to variational quantum circuits. These baselines test the expressivity of pure quantum approaches without classical pre-processing.
Hybrid Baselines.
We include QCL and CQCL, which combine classical convolutional encoders with quantum projection heads or classifiers, representing the current state-of-the-art in hybrid quantum machine learning.
While Transformer-based architectures like Particle Transformer (ParT) (Qu2022ParticleTF) and ParticleNet (qu_jet_2020) achieve state-of-the-art performance (approx. 84% accuracy) on quark-gluon discrimination, they are resource-intensive, utilizing full particle inputs (30–50 constituents) and large parameter spaces (e.g., ParticleNet: $\approx$366k, ParT: $\approx$2.14M). In contrast, this study evaluates QRGCL under a strict NISQ-compatible regime: restricted to 7 input particles and a highly compact total parameter count of $\approx$126k (with only 45 learnable quantum parameters in the rationale generator). Consequently, a direct numerical comparison with full-scale Transformers is not appropriate. We selected GNN baselines as they provide a comparable topological framework to isolate the specific impact of the quantum rationale mechanism within these specific computational constraints.
Table 1: Performance benchmarking and ablation test of the proposed QRGCL. Bolded values indicate the best performance. Results represent the mean and standard deviation calculated over 3 random seeds using the optimal hyperparameter configurations identified in Table 2.
Model Test Acc. ($\uparrow$) AUC ($\uparrow$) F1 Score ($\uparrow$) #Params ($\downarrow$) $n_{\alpha}$ ($\downarrow$) $n_{layer}$ ($\downarrow$) Batch size Epoch Encoder
@@
Parameter breakdown: The reported QRGCL parameter count (126,015) includes the ParticleNet encoder (125,961 params) and the QRG (54 params: 45 quantum + 9 classical), used during pre-training and frozen for fine-tuning. The downstream classifier is a separate linear head (129 params) trained on frozen embeddings.
Table 2: Comparison between classical and quantum RG of RGCL. Reported values represent the mean validation accuracy across stratified 10-fold cross-validation (with shuffling enabled) on the training set. Standard deviations indicate the variability across folds. Bolded values indicate the best performance.
@@
Accuracy ($\uparrow$) AUC ($\uparrow$) F1 score ($\uparrow$) Accuracy ($\uparrow$) AUC ($\uparrow$) F1 score ($\uparrow$)
Nodes per graph 7 $70.4\%\pm 1.2\%$ $\mathbf{76.3\%\pm 1.1\%}$ $69.1\%\pm 1.2\%$ $70.8\%\pm 0.9\%$ $\mathbf{75.9\%\pm 0.8\%}$ $69.9\%\pm 0.9\%$
8 $68.2\%\pm 1.3\%$ $74.1\%\pm 1.2\%$ $67.9\%\pm 1.3\%$ $70.8\%\pm 0.9\%$ $75.6\%\pm 0.9\%$ $70.5\%\pm 0.9\%$
9 $68.4\%\pm 1.4\%$ $73.4\%\pm 1.3\%$ $67.2\%\pm 1.4\%$ $67.6\%\pm 1.1\%$ $74.5\%\pm 1.0\%$ $67.5\%\pm 1.1\%$
10 $67.0\%\pm 1.5\%$ $73.4\%\pm 1.4\%$ $66.5\%\pm 1.5\%$ $70.7\%\pm 0.9\%$ $75.7\%\pm 0.8\%$ $70.3\%\pm 0.9\%$
Number of layers 2 $68.8\%\pm 1.3\%$ $74.5\%\pm 1.2\%$ $8.8\%\pm 1.3\% 66$$6.8\%\pm 1.0\%$ $71.4\%\pm 1.0\%$ $66.5\%\pm 1.0\%$
3 $66.7\%\pm 1.5\%$ $73.6\%\pm 1.4\%$ $66.7\%\pm 1.5\%$ $68.6\%\pm 0.9\%$ $\mathbf{74.7\%\pm 0.8\%}$ $68.5\%\pm 0.9\%$
4 $71.2\%\pm 1.1\%$ $\mathbf{76.3\%\pm 1.0\%}$ $71.1\%\pm 1.1\%$ $66.4\%\pm 1.2\%$ $72.3\%\pm 1.1\%$ $66.4\%\pm 1.2\%$
5 $63.0\%\pm 1.8\%$ $67.5\%\pm 1.7\%$ $63.0\%\pm 1.9\%$ $67.4\%\pm 1.3\%$ $71.2\%\pm 1.2\%$ $67.3\%\pm 1.3\%$
Augmentation ratio 0.0 $65.2\%\pm 1.6\%$ $71.5\%\pm 1.5\%$ $65.1\%\pm 1.7\%$ $69.8\%\pm 1.0\%$ $73.2\%\pm 1.0\%$ $69.5\%\pm 1.0\%$
0.1 $67.8\%\pm 1.4\%$ $\mathbf{74.1\%\pm 1.3\%}$ $67.8\%\pm 1.4\%$ $72.1\%\pm 0.8\%$ $\mathbf{78.8\%\pm 0.7\%}$ $72.1\%\pm 0.8\%$
0.2 $64.6\%\pm 1.7\%$ $69.8\%\pm 1.6\%$ $64.5\%\pm 1.7\%$ $65.6\%\pm 1.4\%$ $71.1\%\pm 1.3\%$ $65.6\%\pm 1.4\%$
0.3 $63.7\%\pm 1.8\%$ $69.6\%\pm 1.7\%$ $63.7\%\pm 1.8\%$ $1.3\%\pm 0.9\% 77$$7.0\%\pm 0.8\%$ $71.3\%\pm 0.9\%$
@@
Table 3: Scalability comparison of classical, quantum, and hybrid models as the number of active nodes per graph ($n_{\alpha}$) increases beyond 3. Entries marked "–" indicate training failures due to out-of-memory (OOM) errors resulting from exponential circuit width growth at $n_{\alpha}=7$. Bolded values indicate the best performance.
Model Test Acc. AUC F1-score $n_{\alpha}$ #Params $n_{\text{layer}}$ Batch size Epoch Encoder
EGNN 54.2% 64.4% 68.5% 7 $52$ 5 64 19 MLP
EQGNN 47.8% 55.1% 30.9% 7 990100 6 1 19 MLP + H
QGNN 54.6% 43.9% 54.6% 7 1021076 6 1 19 MLP + H
GNN 52.2% 57.8% 35.8% 7 $52$ 4 64 19 MLP
QCL 44.8% 48.3% 61.9% 5 384 3 256 1000 Amplitude
QCL 7 3 256 1000 Amplitude
CQCL 45.2% 48.4% 60.7% 5 354 3 128 50 Angle (RY+RX)
CQCL 7 3 128 50 Angle (RY)